OpenStudy (osanseviero):

Manipulating proportions. If we know that a/b = c/d, how can it be (a+c)/(b+d)

OpenStudy (anonymous):

im confused with what you are asking

ganeshie8 (ganeshie8):

do you mean : \[\large \dfrac{a}{b} = \dfrac{c}{d} \implies \dfrac{a+b}{a-b} = \dfrac{c+d}{c-d}\] ?

OpenStudy (osanseviero):

\[\frac{ a }{ b } = \frac{ c }{ d } = \frac{ a + c }{ b + d } = \frac{ a-c }{ b-d }\]

OpenStudy (osanseviero):

I don't understand how does the author do this

ganeshie8 (ganeshie8):

they all are forms of componendo and dividendo rule

ganeshie8 (ganeshie8):

lets see if we can derive this particular rule

OpenStudy (osanseviero):

How can I prove it?

OpenStudy (osanseviero):

I just saw a way to prove from (a+kc)/(b+kd)...is there any way to prove it in a way that is not backwards?

ganeshie8 (ganeshie8):

try adding c/b both sides

OpenStudy (osanseviero):

Which part? The ones of the beginning? I would have (a+c)/b = (c/d)+(c/b)

ganeshie8 (ganeshie8):

yes simplify right hand side

ganeshie8 (ganeshie8):

\[\dfrac{a}{b} = \dfrac{c}{d}\] add \(\large \dfrac{c}{b}\) both sides : \[\large \dfrac{a}{b}+\dfrac{c}{b} = \dfrac{c}{d}+\dfrac{c}{b} \implies \dfrac{a+c}{b+d} = \dfrac{c}{d} \tag{1} \] subtract \(\large \dfrac{c}{b}\) both sides : \[\large \dfrac{a}{b}-\dfrac{c}{b} = \dfrac{c}{d}-\dfrac{c}{b} \implies \dfrac{a-c}{b-d} = \dfrac{c}{d} \tag{2} \]

ganeshie8 (ganeshie8):

the result follolws from (1) & (2)

ganeshie8 (ganeshie8):

Or you could simply apply the regular componendo and dividendo to right hand side

ganeshie8 (ganeshie8):

add +1 to both sides of right hand side and simplify

OpenStudy (osanseviero):

Perfect, thanks a lot :9

ganeshie8 (ganeshie8):

np :) this is really a very useful property http://en.wikipedia.org/wiki/Componendo_and_dividendo

OpenStudy (osanseviero):

:)

OpenStudy (osanseviero):

I will look it deeper later